Formula for the lateral surface area of a cylinder. SA=2(pi)rh, solve for r
Answer by colliefan(242)

SA=2(pi)rh
What you do to one side you must do to the other.
SA*1/2=2*1/2(pi)rh
SA/(2)=(pi)rh
SA/(2*pi)=(pi)*rh*1/(pi)
SA/(2*pi)=rh
SA/(2*pi*h)=rh*1/h
SA/(2*pi*h)=r
r=SA/(2*pi*h)